# ReefNet: A Large-scale, Taxonomically Enriched Dataset and Benchmark for Hard Coral Classification

### v1

@@ 5.3 Zero-shot Results
From Table LABEL:tab:zero_shot, it is evident that all models struggle with the zero-shot classification task, particularly when compared to the fine-tuned models discussed in section 5.2. This performance gap is expected, as most of the zero-shot models have likely encountered limited or no coral-specific data during pretraining. However, BioCLIP (Stevens et al., 2023a), which achieves the best zero-shot performance, has been pre-trained on the TreeOfLife-10M (Stevens et al., 2023c) dataset, which includes a wide range of biological species, including corals. This pretraining provides BioCLIP with prior knowledge of coral-related visual features. Therefore, the performance of VLMs is consistent with their training data exposure.
We also evaluate Qwen2.5-VL (Bai et al., 2025) under different configurations. Directly prompting Qwen2.5-VL to identify the coral genus from an image results in subpar performance. To improve this, we provided additional context by generating coral genus descriptions using GPT-4.0; we refer to this setup as Qwen-GPT. While this improved Qwen2.5-VL’s performance on the in-source split, the performance on the cross-source split remained unchanged. Upon inspection, we found that the GPT-generated descriptions contained repetitive and generic content, limiting their utility. To address this issue, we curated high-quality textual sources by extracting coral-specific information from three reference books (Veron, 2000a; b; Wallace, 1999). These excerpts were then summarized by GPT-4.0 into discriminative genus descriptions, which were provided as context to Qwen2.5-VL; we refer to this setup as Qwen-Book. This approach improved Qwen2.5-VL’s performance on both in-source and cross-source splits. Although Qwen2.5-VL still lags behind BioCLIP, these results highlight the potential of using high-quality, domain-specific textual context to enhance coral recognition in MLLMs. We believe that further refining the descriptions and retrieval strategies could significantly close the gap.
@@
Table 5: Zero-Shot Classification. Macro recall of open-source VLMs on hard corals classification. Qwen-GPT uses GPT-4o-generated genus descriptions; Qwen-Book adds GPT-4o summaries of domain-specific books (Veron, 2000a; b; Wallace, 1999).
@@ 9 Acknowledgements
We gratefully acknowledge the CoralNet team and the many reef researchers who contributed publicly available data to the CoralNet platform, enabling the creation of the ReefNet dataset. We also thank the Red Sea Global team, led by Ivor Williams, Luis Silva, Rhonda Suka, Karla Gonzalez for their support. Special thanks go to Brian Nieuwenhuis, Yesmeen Khattab, Natalie Dunn, Marta Andrea Ezeta Watts, Susanne Bahr, Federica Barreca, Nicolas Oury, and Tullia Isotta Terraneo for their critical efforts in designing the label set, standardizing taxonomy, and mapping labels to the WoRMS (World Register of Marine Species) database and doing the manual verification and quality check on the dataset in general. This research was funded by the Ocean Science and Solutions Applied Research Institute (OSSARI), NEOM Company under the “Towards AI Accelerated Reef Ecology Practice: Expert-informed Benchmarks and Machine Learning Systems for Coral Reef 2D /3D Visual Understanding” project grant (RGC/3/6029-01) awarded to M. Elhoseiny, B. Jones, and F. Benzoni.
@@ A.1 ReefNet in Relation to CoralNet
Despite CoralNet’s (Beijbom et al., 2015) wide adoption as a platform for coral annotations and publishing datasets, it presents several challenges that limit its usability for ML practitioners to train and evaluate ML models in a straightforward way, unlike other biological fields where the community has built well-established ML-ready datasets(e.g., CUB-200-2011, NABirds, iNaturalist, IP102) (Wah et al., 2011; Van Horn et al., 2015; 2018; Wu et al., 2019). ReefNet addresses these shortcomings, making it a more reliable resource for scientific exploration and algorithmic development. The key limitations of CoralNet are:
@@
- World Register of Marine Species (WoRMS) (Board, 2024)-based unified taxonomic labeling: ReefNet adopts a unified labeling scheme grounded in WoRMS to integrate heterogeneous sources. This harmonization also enabled the inclusion of a Red Sea–focused dataset. We anticipate that this taxonomy will encourage future work to adopt scientifically sound annotation practices.
@@ A.3.1 Training Setup
We evaluated a self-supervised variant using a Masked Autoencoder (MAE)He et al. (2022) with a ViT-LargeDosovitskiy et al. (2021) backbone. Pretraining was performed on 1.2 million $512\times 512$ patches of hard coral annotations, using data from an earlier stage of the ReefNet pipeline prior to the final filtration step (i.e., using 85 sources, before removing nine low-quality sources). The MAE was trained for 800 epochs with a 75% masking ratio and cosine learning rate scheduling (base LR $1.5\times 10^{-4}$).
@@
All models were trained using $512\times 512$ patches extracted around point annotations from full-resolution CoralNet imagery, referenced via a custom CSV loader. We selected this patch size to balance local texture and global morphological structure, both of which are essential for accurate coral genus classification. Preliminary experiments with smaller patches ($224\times 224$ and $448\times 448$) showed diminished performance, likely due to loss of spatial context or insufficient detail. This design choice aligns with prior work showing that effective coral classification requires capturing both textural and structural cues Almazán et al. (2019). Each training sample in our data includes genus annotation or family in the case of Fungiidae, patch coordinates, and source metadata.
@@
In our loss function ablation (Table 4), we evaluated several class-balanced and standard variants. We used Cross-Entropy (CE) Loss implemented with PyTorch nn.CrossEntropyLoss(). For Focal Loss, we set $\gamma=2$ and $\alpha=0.25$ (Lin et al., 2017). For the Class-Balanced Cross-Entropy, we applied effective number reweighting with $\beta=0.9999$ (Cui et al., 2019). And for the Class-Balanced Focal Loss, we set $\beta=0.9999$ and $\gamma=2$ (Cui et al., 2019).
@@ A.4.1 Closed-Set Evaluation and Open‑Set Recognition
We implemented a standard open‑set recognition (OSR) protocol similar to  (Vaze et al., 2022; Wang et al., 2024) on the cross‑source split (Train‑S4 / Test‑S3&S4). Specifically, we held out 8 coral classes as unknown (e.g., Agaricia, Pavona), training only on the remaining 30 known classes. At test time, evaluation was performed on a mixed set containing both known (32,173 samples) and unknown (1,867 samples) classes. We followed a stratified hold‑out protocol commonly used in OSR literature by excluding low‑frequency classes and selecting unknowns with $\geq 100$ test‑sample support for sound evaluation. Following (Vaze et al., 2022), we note that a classifier’s closed‑set accuracy correlates strongly with its open‑set detection ability. Accordingly, we use Max Logit Score (MLS) scoring as a competitive baseline, which their experiments show performs on par with specialized OSR methods. Consistent with (Wang et al., 2024), we evaluate multiple scoring rules: 1) Max Softmax Probability (MSP), 2) Max Logit Score (MLS), 3) Energy Score (Liu et al., 2020), and observe in Table 6 that magnitude‑sensitive approaches like MLS and Energy yield the best performance, reinforcing their finding that such scoring functions generalize across both OSR and OOD tasks.
@@ A.4.2 Augmentation Ablation Results
While this work does not introduce new domain adaptation methods, our cross-domain experiments leveraged strong domain adaptation techniques, particularly through diverse data augmentation strategies. Specifically, we used a combination of RandAugment, color jitter, horizontal flipping, Mixup, CutMix, and random erasing (reprob = 0.1). These augmentations are widely used to improve generalization in the presence of domain shift due to variations in lighting, turbidity, and reef structure (Cubuk et al., 2020; Yun et al., 2019). To validate the contribution of these augmentations, we conducted an ablation study on the cross-source benchmark (Train-S4 / Test-S3&S4) using ViT-B (Dosovitskiy et al., 2021). The results in Table 7 demonstrate that using all augmentations yields the highest performance, confirming their collective benefit. Disabling any single augmentation (e.g., RandAug, ColorJitter, Random Erase) results in a performance drop, showing that each plays a role in enhancing robustness. Training without augmentation or with only a single augmentation significantly underperforms, validating that augmentation functions as an effective domain adaptation mechanism in our setup.
@@ A.4.4 Overall Accuracy
EfficientNet Tan & Le (2019) 97.24 96.53 80.79 81.15 93.56 86.90
ResNet He et al. (2016) 93.71 93.53 80.83 80.79 89.74 84.65
ConvNext Liu et al. (2022) 96.89 98.63 80.86 81.08 92.12 83.99
ViT Dosovitskiy et al. (2021) 97.27 98.05 81.08 81.02 93.75 87.95
ViT (scratch) Dosovitskiy et al. (2021) 86.38 89.27 80.72 80.90 92.92 87.51
BEiT Bao et al. (2022) 96.35 96.77 81.46 80.80 92.48 88.03
DeiT Touvron et al. (2021) 96.99 97.20 81.21 81.70 92.72 87.37
Swin Liu et al. (2021) 96.96 96.38 81.23 81.30 92.08 83.19
ViT Dosovitskiy et al. (2021) (MAE He et al. (2022) pretrained) 97.71 96.51 84.05 83.09 94.32 81.74
BioCLIP-FT Stevens et al. (2023b) 89.39 93.54 66.14 70.03 79.40 71.08
@@ A.4.5 Precision and F1 Score
EfficientNet Tan & Le (2019) 76.44 81.41 51.58 47.60 68.30 63.13
ResNet He et al. (2016) 64.96 65.14 44.69 33.21 64.00 54.81
ConvNext Liu et al. (2022) 75.33 80.37 47.81 51.02 67.98 64.12
ViT Dosovitskiy et al. (2021) 73.60 76.93 53.06 49.66 68.65 64.41
ViT (scratch) Dosovitskiy et al. (2021) 68.73 73.41 45.88 49.73 66.21 64.19
BEiT Bao et al. (2022) 71.12 72.14 50.38 37.21 64.77 62.54
DeiT Touvron et al. (2021) 72.68 75.23 50.20 47.14 64.09 64.34
Swin Liu et al. (2021) 75.41 81.39 50.13 48.07 65.56 62.62
ViT Dosovitskiy et al. (2021) (MAE He et al. (2022) pretrained) 77.36 75.33 57.57 51.77 70.22 59.80
BioCLIP-FT Stevens et al. (2023b) 65.16 69.06 66.14 70.03 64.99 59.73
@@
EfficientNet Tan & Le (2019) 74.66 80.48 55.73 55.33 67.32 61.72
ResNet He et al. (2016) 66.04 65.82 54.31 39.73 63.86 58.96
ConvNext Liu et al. (2022) 73.21 78.04 55.11 55.71 66.81 60.61
ViT Dosovitskiy et al. (2021) 71.80 75.00 55.93 55.66 67.62 62.72
ViT (scratch) Dosovitskiy et al. (2021) 63.23 66.57 54.57 55.47 66.25 62.36
BEiT Bao et al. (2022) 70.08 70.70 56.35 44.09 65.29 62.77
DeiT Touvron et al. (2021) 71.08 73.19 56.05 55.57 64.99 62.17
Swin Liu et al. (2021) 73.25 80.17 56.01 55.60 65.49 60.08
ViT Dosovitskiy et al. (2021) (MAE He et al. (2022) pretrained) 74.92 73.10 59.00 57.47 68.76 58.33
BioCLIP-FT Stevens et al. (2023b) 57.43 76.33 79.40 71.08 60.59 56.45
@@ A.6 Text Description Generation and performance for Zero-Shot models
We use two approaches to generate descriptions for various coral genera, both using GPT-4o. 1) We directly prompt GPT-4o to generate a description for a coral genus with the prompt presented in Figure 14. 2) In the second approach, we provide context information about the genus extracted from domain experts’ chosen books Wallace (1999); Veron (2000a; b) and online sources by allowing an agent to scrape results from Google search. The second approach was done using an agentic pipeline built using CrewAI. The system included one agent per book, one agent for online sources, and a final agent to take the information from the previous agents and produce the final output.
@@
From Table 12, it can be seen that across the three vision language models trained on general data from the internet, OpenCLip performs the best. BioCLIP clearly outperformed all the other models across all the splits, supporting the idea that its pretrained data from TreeOfLife-10M Stevens et al. (2023c), which contains coral information, supports its performance.
@@ A.7 Limitations
  • Dynamic Taxonomy: Coral taxonomy is an evolving field of science, and taxonomic classifications are subject to revision over time. While the dataset reflects the most current understanding at the time of compilation, future changes in taxonomy may render some annotations outdated. Users are encouraged to leverage the provided AphiaIDs to verify the latest accepted taxonomy through the World Register of Marine Species (WoRMS) (Board, 2024).
  • Patch-Based Annotations: ReefNet annotations are exclusively patch-based, which may not fully capture the spatial variability and ecological context of coral reefs (e.g., Figure 4). Patch-based models, while valuable, may be less accurate than models leveraging more comprehensive semantic segmentation. Although recent large-scale segmentation datasets, such as CoralVOS and CoralSCOP (Zheng et al., 2024; Ziqiang et al., 2023), offer valuable insights, they lack fine-grained taxonomic data. ReefNet complements these datasets by providing extensive taxonomically detailed annotations, filling a critical gap for future integration with segmentation-based approaches.

### v2

@@ 5.3 Zero-shot Results
From Table LABEL:tab:zero_shot, it is evident that all models struggle with the zero-shot classification task, particularly when compared to the fine-tuned models discussed in section 5.2. This performance gap is expected, as most of the zero-shot models have likely encountered limited or no coral-specific data during pretraining. However, BioCLIP (BioCLIP), which achieves the best zero-shot performance, has been pre-trained on the TreeOfLife-10M (treeoflife_10m) dataset, which includes a wide range of biological species, including corals. This pretraining provides BioCLIP with prior knowledge of coral-related visual features. Therefore, the performance of VLMs is consistent with their training data exposure.
We also evaluate Qwen2.5-VL (qwen25) under different configurations. Directly prompting Qwen2.5-VL to identify the coral genus from an image results in subpar performance. To improve this, we provided additional context by generating coral genus descriptions using GPT-4.0; we refer to this setup as Qwen-GPT. While this improved Qwen2.5-VL’s performance on the in-source split, the performance on the cross-source split remained unchanged. Upon inspection, we found that the GPT-generated descriptions contained repetitive and generic content, limiting their utility. To address this issue, we curated high-quality textual sources by extracting coral-specific information from three reference books (coralvol1; coralvol3; staghorn). These excerpts were then summarized by GPT-4.0 into discriminative genus descriptions, which were provided as context to Qwen2.5-VL; we refer to this setup as Qwen-Book. This approach improved Qwen2.5-VL’s performance on both in-source and cross-source splits. Although Qwen2.5-VL still lags behind BioCLIP, these results highlight the potential of using high-quality, domain-specific textual context to enhance coral recognition in MLLMs. We believe that further refining the descriptions and retrieval strategies could significantly close the gap.
@@
Table 5: Zero-Shot Classification. Macro recall of open-source VLMs on hard corals classification. Qwen-GPT uses GPT-4o-generated genus descriptions; Qwen-Book adds GPT-4o summaries of domain-specific books (coralvol1; coralvol3; staghorn).
@@ 9 Acknowledgements
We gratefully acknowledge the CoralNet team and the many reef researchers who contributed publicly available data to the CoralNet platform, enabling the creation of the ReefNet dataset. We also thank the Red Sea Global team, led by Ivor Williams, Luis Silva, Rhonda Suka, Karla Gonzalez for their support. Special thanks go to Brian Nieuwenhuis, Yesmeen Khattab, Natalie Dunn, Marta Andrea Ezeta Watts, Susanne Bahr, Federica Barreca, Nicolas Oury, and Tullia Isotta Terraneo for their critical efforts in designing the label set, standardizing taxonomy, and mapping labels to the WoRMS (World Register of Marine Species) database and doing the manual verification and quality check on the dataset in general. This research was funded by the Ocean Science and Solutions Applied Research Institute (OSSARI), NEOM Company under the “Towards AI Accelerated Reef Ecology Practice: Expert-informed Benchmarks and Machine Learning Systems for Coral Reef 2D /3D Visual Understanding” project grant (RGC/3/6029-01) awarded to M. Elhoseiny, B. Jones, and F. Benzoni. For computer time, this research used Ibex managed by the Supercomputing Core Laboratory at King Abdullah University of Science & Technology (KAUST) in Thuwal, Saudi Arabia.
@@ A.1 ReefNet in Relation to CoralNet
Despite CoralNet’s (beijbom2015coralnet) wide adoption as a platform for coral annotations and publishing datasets, it presents several challenges that limit its usability for ML practitioners to train and evaluate ML models in a straightforward way, unlike other biological fields where the community has built well-established ML-ready datasets(e.g., CUB-200-2011, NABirds, iNaturalist, IP102) (wah2011cub; vanhorn2015nabirds; vanhorn2018inaturalist; wu2019ip102). ReefNet addresses these shortcomings, making it a more reliable resource for scientific exploration and algorithmic development. The key limitations of CoralNet are:
@@
- World Register of Marine Species (WoRMS) (worms)-based unified taxonomic labeling: ReefNet adopts a unified labeling scheme grounded in WoRMS to integrate heterogeneous sources. This harmonization also enabled the inclusion of a Red Sea–focused dataset. We anticipate that this taxonomy will encourage future work to adopt scientifically sound annotation practices.
@@ A.3.1 Training Setup
We evaluated a self-supervised variant using a Masked Autoencoder (MAE)he2022masked with a ViT-Largevit backbone. Pretraining was performed on 1.2 million $512\times 512$ patches of hard coral annotations, using data from an earlier stage of the ReefNet pipeline prior to the final filtration step (i.e., using 85 sources, before removing nine low-quality sources). The MAE was trained for 800 epochs with a 75% masking ratio and cosine learning rate scheduling (base LR $1.5\times 10^{-4}$).
@@
All models were trained using $512\times 512$ patches extracted around point annotations from full-resolution CoralNet imagery, referenced via a custom CSV loader. We selected this patch size to balance local texture and global morphological structure, both of which are essential for accurate coral genus classification. Preliminary experiments with smaller patches ($224\times 224$ and $448\times 448$) showed diminished performance, likely due to loss of spatial context or insufficient detail. This design choice aligns with prior work showing that effective coral classification requires capturing both textural and structural cues almazan2019coral. Each training sample in our data includes genus annotation or family in the case of Fungiidae, patch coordinates, and source metadata.
@@
In our loss function ablation (Table 4), we evaluated several class-balanced and standard variants. We used Cross-Entropy (CE) Loss implemented with PyTorch nn.CrossEntropyLoss(). For Focal Loss, we set $\gamma=2$ and $\alpha=0.25$ (focal_loss). For the Class-Balanced Cross-Entropy, we applied effective number reweighting with $\beta=0.9999$ (Cui2019ClassBalancedLB). And for the Class-Balanced Focal Loss, we set $\beta=0.9999$ and $\gamma=2$ (Cui2019ClassBalancedLB).
@@ A.4.1 Closed-Set Evaluation and Open‑Set Recognition
We implemented a standard open‑set recognition (OSR) protocol similar to  (vaze2022openset; wang2024dissect) on the cross‑source split (Train‑S4 / Test‑S3&S4). Specifically, we held out 8 coral classes as unknown (e.g., Agaricia, Pavona), training only on the remaining 30 known classes. At test time, evaluation was performed on a mixed set containing both known (32,173 samples) and unknown (1,867 samples) classes. We followed a stratified hold‑out protocol commonly used in OSR literature by excluding low‑frequency classes and selecting unknowns with $\geq 100$ test‑sample support for sound evaluation. Following (vaze2022openset), we note that a classifier’s closed‑set accuracy correlates strongly with its open‑set detection ability. Accordingly, we use Max Logit Score (MLS) scoring as a competitive baseline, which their experiments show performs on par with specialized OSR methods. Consistent with (wang2024dissect), we evaluate multiple scoring rules: 1) Max Softmax Probability (MSP), 2) Max Logit Score (MLS), 3) Energy Score (NEURIPS2020_f5496252), and observe in Table 6 that magnitude‑sensitive approaches like MLS and Energy yield the best performance, reinforcing their finding that such scoring functions generalize across both OSR and OOD tasks.
@@ A.4.2 Augmentation Ablation Results
While this work does not introduce new domain adaptation methods, our cross-domain experiments leveraged strong domain adaptation techniques, particularly through diverse data augmentation strategies. Specifically, we used a combination of RandAugment, color jitter, horizontal flipping, Mixup, CutMix, and random erasing (reprob = 0.1). These augmentations are widely used to improve generalization in the presence of domain shift due to variations in lighting, turbidity, and reef structure (randaugment; yun2019cutmix). To validate the contribution of these augmentations, we conducted an ablation study on the cross-source benchmark (Train-S4 / Test-S3&S4) using ViT-B (vit). The results in Table 7 demonstrate that using all augmentations yields the highest performance, confirming their collective benefit. Disabling any single augmentation (e.g., RandAug, ColorJitter, Random Erase) results in a performance drop, showing that each plays a role in enhancing robustness. Training without augmentation or with only a single augmentation significantly underperforms, validating that augmentation functions as an effective domain adaptation mechanism in our setup.
@@ A.4.4 Overall Accuracy
EfficientNet tan2019efficientnet 97.24 96.53 80.79 81.15 93.56 86.90
ResNet he2016deep 93.71 93.53 80.83 80.79 89.74 84.65
ConvNext liu2022convnet 96.89 98.63 80.86 81.08 92.12 83.99
ViT vit 97.27 98.05 81.08 81.02 93.75 87.95
ViT (scratch) vit 86.38 89.27 80.72 80.90 92.92 87.51
BEiT bao2021beit 96.35 96.77 81.46 80.80 92.48 88.03
DeiT deit 96.99 97.20 81.21 81.70 92.72 87.37
Swin swin 96.96 96.38 81.23 81.30 92.08 83.19
ViT vit (MAE he2022masked pretrained) 97.71 96.51 84.05 83.09 94.32 81.74
BioCLIP-FT bioclip2023 89.39 93.54 66.14 70.03 79.40 71.08
@@ A.4.5 Precision and F1 Score
EfficientNet tan2019efficientnet 76.44 81.41 51.58 47.60 68.30 63.13
ResNet he2016deep 64.96 65.14 44.69 33.21 64.00 54.81
ConvNext liu2022convnet 75.33 80.37 47.81 51.02 67.98 64.12
ViT vit 73.60 76.93 53.06 49.66 68.65 64.41
ViT (scratch) vit 68.73 73.41 45.88 49.73 66.21 64.19
BEiT bao2021beit 71.12 72.14 50.38 37.21 64.77 62.54
DeiT deit 72.68 75.23 50.20 47.14 64.09 64.34
Swin swin 75.41 81.39 50.13 48.07 65.56 62.62
ViT vit (MAE he2022masked pretrained) 77.36 75.33 57.57 51.77 70.22 59.80
BioCLIP-FT bioclip2023 65.16 69.06 66.14 70.03 64.99 59.73
@@
EfficientNet tan2019efficientnet 74.66 80.48 55.73 55.33 67.32 61.72
ResNet he2016deep 66.04 65.82 54.31 39.73 63.86 58.96
ConvNext liu2022convnet 73.21 78.04 55.11 55.71 66.81 60.61
ViT vit 71.80 75.00 55.93 55.66 67.62 62.72
ViT (scratch) vit 63.23 66.57 54.57 55.47 66.25 62.36
BEiT bao2021beit 70.08 70.70 56.35 44.09 65.29 62.77
DeiT deit 71.08 73.19 56.05 55.57 64.99 62.17
Swin swin 73.25 80.17 56.01 55.60 65.49 60.08
ViT vit (MAE he2022masked pretrained) 74.92 73.10 59.00 57.47 68.76 58.33
BioCLIP-FT bioclip2023 57.43 76.33 79.40 71.08 60.59 56.45
@@ A.6 Text Description Generation and performance for Zero-Shot models
We use two approaches to generate descriptions for various coral genera, both using GPT-4o. 1) We directly prompt GPT-4o to generate a description for a coral genus with the prompt presented in Figure 14. 2) In the second approach, we provide context information about the genus extracted from domain experts’ chosen books staghorn; coralvol1; coralvol3 and online sources by allowing an agent to scrape results from Google search. The second approach was done using an agentic pipeline built using CrewAI. The system included one agent per book, one agent for online sources, and a final agent to take the information from the previous agents and produce the final output.
@@
From Table 12, it can be seen that across the three vision language models trained on general data from the internet, OpenCLip performs the best. BioCLIP clearly outperformed all the other models across all the splits, supporting the idea that its pretrained data from TreeOfLife-10M treeoflife_10m, which contains coral information, supports its performance.
@@ A.7 Limitations
  • Dynamic Taxonomy: Coral taxonomy is an evolving field of science, and taxonomic classifications are subject to revision over time. While the dataset reflects the most current understanding at the time of compilation, future changes in taxonomy may render some annotations outdated. Users are encouraged to leverage the provided AphiaIDs to verify the latest accepted taxonomy through the World Register of Marine Species (WoRMS) (worms).
  • Patch-Based Annotations: ReefNet annotations are exclusively patch-based, which may not fully capture the spatial variability and ecological context of coral reefs (e.g., Figure 4). Patch-based models, while valuable, may be less accurate than models leveraging more comprehensive semantic segmentation. Although recent large-scale segmentation datasets, such as CoralVOS and CoralSCOP (coralscop; coralvos), offer valuable insights, they lack fine-grained taxonomic data. ReefNet complements these datasets by providing extensive taxonomically detailed annotations, filling a critical gap for future integration with segmentation-based approaches.